# Superposition in Graph Neural Networks

### v1

@@ 1 Introduction
  • Width produces a phase pattern. As the final hidden layer widens, feature overlap first decreases, then briefly increases around the point where capacity matches the number of concepts, then decreases again. At high width, graph-level features approach near-ideal packing, but node-level concepts often remain entangled – evidence that readout layers can disentangle what the message-passing stack keeps mixed.
  • Pooling sharpness encourages axis alignment. Making pooling more "winner-take-all" (e.g., max pooling) increases alignment to coordinate axes and reduces node-level feature sharing. We explain this with two complementary arguments: gradients concentrate on large coordinates, and under noise the axis-aligned choice loses less information than an oblique one.
  • Rank collapse can appear even in shallow GNNs. In some runs the numerical rank of the pooled representation stays below the number of learned features while the accuracy remains high. We link this to hard gating after aggregation (exactly zeroing channels) and to a loss-driven preference for "mutually obtuse" class directions that resists activating new dimensions.
Contributions. (i) A representation-centric framework for analyzing superposition in GNNs that works across architectures and layers; (ii) robust, simple diagnostics and feature extraction procedures (centroids and probes) that avoid assumptions about neurons or coordinates; (iii) empirical maps of how width, topology, and pooling shape superposition; and (iv) practical guidance – when to expect entanglement, when pooling helps, and how to avoid low-rank traps.
@@ 2.1 Superposition and the interpretability of internal representations
Classic unit analyses in CNNs found channels that act like concept detectors (Zeiler and Fergus, 2014; Zhou et al., 2014; Bau et al., 2017; Olah et al., 2017); related work spans RNNs and transformers (Karpathy et al., 2015; Strobelt et al., 2017; Gurnee et al., 2023, 2024; Geva et al., 2020), and early GNN studies exist (Xuanyuan et al., 2023). However, concepts need not be axis-aligned: they can correspond to arbitrary directions or nonlinear detectors. Directional tools (TCAV, linear probes) therefore treat a feature as a vector in latent space (Kim et al., 2018; Alain and Bengio, 2016); recent sparse-autoencoder work aims to demix polysemantic units (Bricken et al., 2023; Gao et al., 2024).
@@ 2.2 Representations in GNNs
A message-passing layer aggregates neighbor information and applies a local transform,
@@
with architecture-specific aggregation (e.g., normalized sum for GCN (Kipf and Welling, 2017), unnormalized sum + MLP for GIN (Xu et al., 2019), attention for GAT Veličković et al. (2018)). After $L$ layers, node embeddings $\mathbf{H}^{(L)}$ are pooled into a graph embedding $\mathbf{h}_{G}=g(\mathbf{H}^{(L)})$ via sum/mean/max.
@@ 3.1 Interpreting and extracting features
$\mathbf{X}\xrightarrow{\phi_{1}}\mathbf{H}^{(1)}\xrightarrow{\phi_{2}}\cdots\xrightarrow{\phi_{L}}\mathbf{H}^{(L)}\rightarrow\mathbf{h}_{G}\xrightarrow{g}\hat{\mathbf{y}}$,
each $\phi_{\ell}$ is a message-passing block, $\mathbf{H}^{(\ell)}\!\in\!\mathbb{R}^{n_{\ell}\times d_{\ell}}$ are node embeddings, $g$ is a graph-pool, and $\mathbf{h}_{G}\!\in\!\mathbb{R}^{d}$ is the graph embedding. We interpret features in two complementary ways.
(1) Linear-probe features (model-decodable concepts).
Given embeddings $\{\mathbf{h}_{G}\}$ (graph-level) or rows of $\mathbf{H}^{(\ell)}$ (node-level) and binary targets $y_{\ell}\!\in\!\{0,1\}$ for concept $\ell$, we fit a logistic probe on a held-out split:
$s_{\ell}=\mathbf{w}_{\ell}^{\top}\mathbf{z}+b_{\ell},\quad\mathbf{z}\in\{\mathbf{h}_{G}\}\ \text{or}\ \{\mathbf{H}^{(\ell)}_{i:}\}$.
@@
(2) Class-conditional centroids (task-aligned concepts). On a held-out split, we form ground-truth one-hot sets
@@
and stack rows to obtain $C\in\mathbb{R}^{K\times d}$. In the linear/LDA regime, discriminants align with class means (Alain and Bengio, 2016), so centroid directions approximate task-discriminative axes and aggregate the combined effects of message passing and pooling.
@@
  • Centroids: class $\ell$ is active if in-class recall $\geq 0.5$ and all off-diagonal recalls <0.5<0.5.
@@ 3.2 Proposed Metrics
With singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots$ and $p_{i}=\sigma_{i}/\sum_{j}\sigma_{j}$, the entropy-based effective rank is
@@
It estimates the number of effectively used axes. For centroids we COM-center $C^{\circ}=C-\mathbf{1}\bar{\mathbf{c}}^{\top}$ with $\bar{\mathbf{c}}=\tfrac{1}{k_{a}}\sum_{\ell}\mathbf{c}_{\ell}$ before computing EffRank, removing a global bias direction that no practitioner would interpret as a mechanism. For probe normals we use the raw $C$.
@@
Welch-Normalized Overlap (WNO). While SI measures the mean number of features distributed over each axis, it does not consider their angular geometry [^1] For example, two different arrangements of packing 6 vectors into 2D: (i) arranged as a near-regular hexagon, and (ii) arranged such that there are two antipodal clusters. Both have $\text{EffRank}\approx 2$.. Therefore to complement SI, we introduce an angular overlap metric called the Welch-normalized overlap: let $\widetilde{C}$ be a row-unit–normalized projection of $C$ and define
@@
so $\mathrm{WNO}=0$ is Welch-optimal (least overlap), $\mathrm{WNO}=1$ is random, and $\mathrm{WNO}>1$ is worse-than-random packing. Since ambient $d$ can be much larger than the subspace actually used. We therefore report intrinsic WNO: set $r=\lceil\mathrm{EffRank}(C^{\dagger})\rceil$ (with $C^{\dagger}=C^{\circ}$ for centroids and $C^{\dagger}=C$ for probes), project $C$ to the top-$r$ right-singular subspace to obtain $C_{r}$, row-normalize, and take $d_{\mathrm{eff}}=r$ in (4). For centroids and graph-level probes we additionally remove a single shared offset direction before the SVD (PC1 removal); for node-level probes we do not remove PC1.222PC1 removal is appropriate when a single global bias dominates all classes (e.g., pooled magnitude); removing it improves sensitivity to relative packing. For node probes this shared offset is not guaranteed and we keep the raw geometry. If $r\leq 1$ or $k_{a}\leq 1$, WNO is undefined and we report NA.
@@
EffRank and WNO are invariant to right-multiplication by orthogonal matrices (basis changes) and to uniform rescaling of rows; they probe geometry intrinsic to the representation rather than coordinate choices. SI quantifies features-per-axis pressure; WNO quantifies how tightly those axes are packed relative to principled baselines.
@@ 4.1 Datasets
The Pairwise dataset. We introduce a dataset where graphs are equal length linear chains. Each node has an $n$-dimensional one-hot/zero feature. Class $k$ is active if two adjacent nodes both activate dimension $k$, yielding sparse multi-class targets. The task is attribute driven: since each class is associated with identical graph structure, we can probe for superposition whilst factoring out the effects of varying topology. Sparsity is controlled by the activation probability $p$ of a node having a one-hot (non-zero) feature.
@@
All other graphs receive $y_{A}=y_{B}=0$. Thus the same lower-level motif concepts $\{C_{3},C_{4},C_{5},C_{6}\}$ are re-used across the two tasks, but each label requires a distinct pair. This cleanly separates (i) the formation of topological features from (ii) how a model combines them. A graph is formed by sampling required counts for the four motifs according to the label pattern and then optionally adding extra copies of each motif with independent probability $p_{\text{extra}}$. Motif instances are placed on disjoint node sets and connected by random bridges to avoid trivial overlaps. Because there are no informative node attributes, any successful solution must construct the motif detectors purely via message passing.
@@ 4.2 Width-induced superposition
We study how width $d$ affects superposition on PAIRWISE ($k=16$). The first hidden layer is set to the input width ($k$) and the second layer to $d$, inducing a bottleneck when $d<k$. We measure (a) class-conditional centroids in the graph space and (b) node-level concepts $\mathsf{Is}_{t}$ and $\mathsf{NextTo}_{t}$ via linear probes on the last pre-pooling layer: let $x_{i}\in\{e_{1},\ldots,e_{k}\}$ be the one-hot node type and $\mathcal{N}_{i}$ the neighbors of $i$. For type $t$,
$\mathsf{Is}_{t}(i):=\{x_{i}=e_{t}\},\qquad\mathsf{NextTo}_{t}(i):=\mathbf{1}\{\exists j\in\mathcal{N}_{i}:\,x_{j}=e_{t}\}.$
We keep only active features: centroids with in-class recall >0.5>0.5 and off-class recall <0.5<0.5, and probes with AUROC >0.6>0.6. We report intrinsic WNO (computed in the EffRank-selected subspace). Centroids are COM-centered for EffRank; probe normals are not.
@@
Figure 1: Superposition effects across bottleneck dimensions $d$. A dashed line corresponds to $k=d$.
@@
Architecture trends follow inductive bias. For GIN, the post-aggregation MLP tends to funnel many labels into shared axes ("generic neighbor/type pressure + small residuals"), which the readout later separates by sign/magnitude at the graph level. GAT shows peaks that are present but typically smaller; attention can separate channels earlier, while GCN produces smoother, less volatile curves.
The WNO/SI peak typically coincides with saturation of the number of active features/probes and the onset of a flat test loss. This aligns with two competing effects of more width: feature formation (easier to instantiate additional task-useful directions) versus packing efficiency (easier to spread existing directions). Early on, packing dominates; around $d\!\approx\!k$ capacity is spent on forming new directions (SI/WNO rise); once formation saturates, packing dominates again and SI/WNO drop. Centroids are tied directly to the BCE objective, so their feature formation saturates earlier, explaining the shorter first phase.
4.3 Topology-induced superposition
We train 3-layer GCN, GIN, and GATv2 models (16 channels per layer), apply a global $\mathsf{mean}$ pool followed by a ReLU, and fit a two-logit readout $(y_{A},y_{B})$. For each architecture we repeat training over 250 seeds. We study two feature families: (i) node-level probes $\mathsf{Inside}(C_{\ell})$ on the last pre-pooling layer, and (ii) graph-level probes $\mathsf{Has}(C_{\ell})$ on the post-pooling representation.
@@ 4.3 Topology-induced superposition
Despite $d\!\gg\!k$, superposition remains large. With k=4k{=}4 motifs and width d=16d{=}16, independent axes would yield $\text{SI}\!\approx\!1$ and $\text{WNO}\!\approx\!0$. Instead, Fig. 2 shows clusters at $\text{SI}\!>\!1$ and $\text{WNO}\!\gg\!0$, i.e., above random overlap. In most runs, at least one motif pair has $|\cos|>0.9$ (Table 1), most often $(C_{3},C_{4})$ or $(C_{5},C_{6})$; in rare cases ($\sim$1% in GIN) all four are nearly collinear, consistent with an emergent lever of ’cycleness’. High accuracy can still be achieved, suggesting that the pooling / reading takes advantage of residual multicoordinate structure or magnitude even when node-level directions coalesce.
Cosine geometry reveals a cyclic structure. Figure 3 reports mean cosine similarity matrices. At the node level, the lengths of the nearby cycles align (for example, $C_{3}$ with $C_{4}$, $C_{5}$ with $C_{6}$) while distant pairs are weakly aligned or slightly antialigned. An MPNN layer $L$ implements a low-degree polynomial in the normalized adjacency, so detectors for $C_{3}\!\ldots C_{6}$ reuse monomials of closed walks up to length $\approx 2L{+}1$, naturally making adjacent lengths more similar. After pooling, geometry re-mixes toward the task: the post-pool ReLU and linear readout learn graph-level axes that are linear combinations of node-level detectors (e.g., in GIN with mean-pool, $C_{3}$ aligns with $C_{6}$ while $C_{4}$ opposes $C_{5}$, which helps reject single-motif false positives).
Figure 3: Mean cosine similarities for the node features (top) and graph features (bottom).
@@
Max pooling reduces lever sharing. Since a linear probe’s sign can flip with an equivalent boundary, we track $|\cos|$ to detect shared levers. Switching from mean to max pooling reduces node-level alignment across all architectures: both the fraction of runs with any $|\cos|>0.9$ pair and with all pairs >0.9>0.9 drop (Table 1). Intuitively, a per-channel max cannot encode a conjunction on a single channel – one large activation would mask the other—so the model is pressured to place motifs on distinct channels, lowering $|\cos|$. The ordering in Table 1 (GCN >> GAT >> GIN in lever sharing) aligns with each layer’s mixing bias. GCN’s degree normalized sum acts as a low-pass smoother, concentrating signals into a few graph-harmonic axes and driving different motifs to share directions. GAT’s learned attention reduces indiscriminate mixing and yields moderate de-alignment, but a single head offers limited selectivity. GIN’s unnormalized sum plus post aggregation MLP provides the most expressive channel rotation and gating, producing more specialized directions and the lowest node and graph level alignment.
Discussion. Even with generous width, semantically different topology features occupy similar node-space directions. Max pooling alleviates but does not eliminate this, probably reflecting message-passing inductive biases. Note that many high-accuracy models sit at $\text{SI}>1$ and $\text{WNO}\gg 0$ (Fig. 2), indicating that accurate solutions often rely on node-level lever sharing with disentanglement deferred to the graph-level readout – superposition here is not merely a symptom of under-capacity.

### v2

@@ 1 Introduction
  • Width produces a phase pattern. As the final hidden layer widens, feature overlap first decreases, then briefly increases around the point where capacity matches the number of concepts, then decreases again. At high width, graph‑level features approach near‑ideal packing, but node‑level concepts often remain entangled – evidence that readout layers can disentangle what the message‑passing stack keeps mixed.
  • Pooling sharpness encourages axis alignment. Making pooling more "winner‑take‑all" (e.g., max pooling) increases alignment to coordinate axes and reduces node‑level feature sharing. We explain this with two complementary arguments: gradients concentrate on large coordinates, and under noise the axis‑aligned choice loses less information than an oblique one.
  • Rank collapse can appear even in shallow GNNs. In some runs the numerical rank of the pooled representation stays below the number of learned features while the accuracy remains high. We link this to hard gating after aggregation (exactly zeroing channels) and to a loss‑driven preference for "mutually obtuse" class directions that resists activating new dimensions.
Contributions. (i) A representation‑centric framework for analyzing superposition in GNNs that works across architectures and layers; (ii) robust, simple diagnostics and feature extraction procedures (centroids and probes) that avoid assumptions about neurons or coordinates; (iii) empirical maps of how width, topology, and pooling shape superposition; and (iv) practical guidance – when to expect entanglement, when pooling helps, and how to avoid low‑rank traps.
@@ 2.1 Superposition and the interpretability of internal representations
Classic unit analyses in CNNs found channels that act like concept detectors (Zeiler and Fergus, 2014; Zhou et al., 2014; Bau et al., 2017; Olah et al., 2017); related work spans RNNs and transformers (Karpathy et al., 2015; Strobelt et al., 2017; Gurnee et al., 2023, 2024; Geva et al., 2020), and early GNN studies exist (Xuanyuan et al., 2023). However, concepts need not be axis‑aligned: they can correspond to arbitrary directions or nonlinear detectors. Directional tools (TCAV, linear probes) therefore treat a feature as a vector in latent space (Kim et al., 2018; Alain and Bengio, 2016); recent sparse‑autoencoder work aims to demix polysemantic units (Bricken et al., 2023; Gao et al., 2024).
@@ 2.2 Representations in GNNs
A message‑passing layer aggregates neighbor information and applies a local transform,
@@
with architecture‑specific aggregation (e.g., normalized sum for GCN (Kipf and Welling, 2017), unnormalized sum + MLP for GIN (Xu et al., 2019), attention for GAT veličković2018graphattentionnetworks). After $L$ layers, node embeddings $\mathbf{H}^{(L)}$ are pooled into a graph embedding $\mathbf{h}_{G}=g(\mathbf{H}^{(L)})$ via sum/mean/max.
@@ 3.1 Interpreting and extracting features
$\mathbf{X}\xrightarrow{\phi_{1}}\mathbf{H}^{(1)}\xrightarrow{\phi_{2}}\cdots\xrightarrow{\phi_{L}}\mathbf{H}^{(L)}\xrightarrow{g}\mathbf{h}_{G}\rightarrow\hat{\mathbf{y}}$,
each $\phi_{l}$ is a message‑passing block, $\mathbf{H}^{(l)}\!\in\!\mathbb{R}^{n_{l}\times d_{l}}$ are node embeddings, $g$ is a graph‑pool, and $\mathbf{h}_{G}\!\in\!\mathbb{R}^{d}$ is the graph embedding. We interpret features in two complementary ways.
(1) Linear‑probe features (model‑decodable concepts).
Linear probes are a well established method for finding feature directions (Akhondzadeh et al., 2023). Given embeddings $\{\mathbf{h}_{G}\}$ (graph‑level) or rows of $\mathbf{H}^{(l)}$ (node‑level) and binary targets $y_{\ell}\!\in\!\{0,1\}$ for concept $\ell$, we fit a logistic probe on a held‑out split:
$s_{\ell}=\mathbf{w}_{\ell}^{\top}\mathbf{z}+b_{\ell},\quad\mathbf{z}\in\{\mathbf{h}_{G}\}\ \text{or}\ \{\mathbf{H}^{(l)}_{i:}\}$.
@@
(2) Class-conditional centroids (task-aligned concepts). On a held‑out split, we form one‑hot sets using the ground truth $\mathbf{y}$.
@@
and stack rows to obtain $C\in\mathbb{R}^{K\times d}$. In the linear/LDA regime, discriminants align with class means (Alain and Bengio, 2016), so centroid directions approximate task‑discriminative axes and aggregate the combined effects of message passing and pooling.
@@
  • Centroids: class $\ell$ is active if in‑class recall $\geq 0.5$ and all off‑diagonal recalls <0.5<0.5.
@@ 3.2 Proposed Metrics
With singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots$ and $p_{i}=\sigma_{i}/\sum_{j}\sigma_{j}$, the entropy‑based effective rank is
@@
It estimates the number of effectively used axes. For centroids we COM‑center $C^{\circ}=C-\mathbf{1}\bar{\mathbf{c}}^{\top}$ with $\bar{\mathbf{c}}=\tfrac{1}{k_{a}}\sum_{\ell}\mathbf{c}_{\ell}$ before computing EffRank, removing a global bias direction that no practitioner would interpret as a mechanism. For probe normals we use the raw $C$.
@@
Welch-Normalized Overlap (WNO). While SI measures the mean number of features distributed over each axis, it does not consider their angular geometry [^1] For example, two different arrangements of packing 6 vectors into 2D: (i) arranged as a near-regular hexagon, and (ii) arranged such that there are two antipodal clusters. Both have $\text{EffRank}\approx 2$.. Therefore to complement SI, we introduce an angular overlap metric called the Welch-normalized overlap: let $\widetilde{C}$ be the matrix obtained from $C$ by normalizing each row to unit $\ell_{2}$ norm and define
@@
so $\mathrm{WNO}=0$ is Welch‑optimal (least overlap), $\mathrm{WNO}=1$ is random, and $\mathrm{WNO}>1$ is worse‑than‑random packing. Since ambient $d$ can be much larger than the subspace actually used. We therefore report intrinsic WNO: set $r=\lceil\mathrm{EffRank}(C^{\dagger})\rceil$ (with $C^{\dagger}=C^{\circ}$ for centroids and $C^{\dagger}=C$ for probes), project $C$ to the top‑$r$ right‑singular subspace to obtain $C_{r}$, row‑normalize, and take $d_{\mathrm{eff}}=r$ in (4). For centroids and graph‑level probes we additionally remove a single shared offset direction before the SVD (PC1 removal); for node‑level probes we do not remove PC1.222PC1 removal is appropriate when a single global bias dominates all classes (e.g., pooled magnitude); removing it improves sensitivity to relative packing. For node probes this shared offset is not guaranteed and we keep the raw geometry. If $r\leq 1$ or $k_{a}\leq 1$, WNO is undefined and we report NA.
@@
EffRank and WNO are invariant to right‑multiplication by orthogonal matrices (basis changes) and to uniform rescaling of rows; they probe geometry intrinsic to the representation rather than coordinate choices. SI quantifies features‑per‑axis pressure; WNO quantifies how tightly those axes are packed relative to principled baselines.
@@ 4.1 Datasets
The Pairwise dataset. We introduce a dataset where graphs are equal length linear chains of 20 nodes. Each node has an $n$-dimensional one-hot/zero feature. Class $k$ is active if two adjacent nodes both activate dimension $k$, yielding sparse multi-class targets. The task is attribute driven: since each class is associated with identical graph structure, we can probe for superposition whilst factoring out the effects of varying topology. Sparsity is controlled by the activation probability $p$ of a node having a one-hot (non-zero) feature. Exact details for reproduction are included in Appendix A.1.
@@
All other graphs receive $y_{A}=y_{B}=0$. Thus the same lower‑level motif concepts $\{C_{3},C_{4},C_{5},C_{6}\}$ are re‑used across the two tasks, but each label requires a distinct pair. This cleanly separates (i) the formation of topological features from (ii) how a model combines them. A graph is formed by sampling required counts for the four motifs according to the label pattern and then adding $\text{Binomial}(8,0.5)$ extra copies of each motif. Exact details for reproduction are included in Appendix A.2. Motif instances are placed on disjoint node sets and connected by random bridges to avoid trivial overlaps. Because there are no informative node attributes, any successful solution must construct the motif detectors purely via message passing.
@@ 4.2 Width-induced superposition
We study how width $d$ affects superposition on PAIRWISE ($k=16$) on $16$ seeds. The first hidden layer is set to the input width ($k$) and the second layer to $d$, inducing a bottleneck when $d<k$. We measure (a) class‑conditional centroids in the graph space and (b) node‑level concepts $\mathsf{Is}_{t}$ and $\mathsf{NextTo}_{t}$ via linear probes on the last pre‑pooling layer: let $x_{i}\in\{e_{1},\ldots,e_{k}\}$ be the one‑hot node type and $\mathcal{N}_{i}$ the neighbors of $i$. For type $t$,
$\mathsf{Is}_{t}(i):=\mathbf{1}\{x_{i}=e_{t}\},\qquad\mathsf{NextTo}_{t}(i):=\mathbf{1}\{\exists j\in\mathcal{N}_{i}:\,x_{j}=e_{t}\}.$
We keep only active features: centroids with in‑class recall >0.5>0.5 and off‑class recall <0.5<0.5, and probes with AUROC >0.6>0.6. We report intrinsic WNO (computed in the EffRank‑selected subspace). Centroids are COM‑centered for EffRank; probe normals are not. Optimizer and training details for this experiment and subsequent ones are given in B.1.
@@
Figure 1: Superposition effects across bottleneck dimensions $d$. A dashed line corresponds to $k=d$. Shaded regions show uncertainty as mean $\pm 1.96\sigma/\sqrt{R}$ where $R=16$ is the number of seeds.
@@
Architecture trends follow inductive bias. For GIN, the post‑aggregation MLP tends to funnel many labels into shared axes ("generic neighbor/type pressure + small residuals"), which the readout later separates by sign/magnitude at the graph level. GAT shows peaks that are present but typically smaller; attention can separate channels earlier, while GCN produces smoother, less volatile curves.
The WNO/SI peak typically coincides with saturation of the number of active features/probes and the onset of a flat test loss. This aligns with two competing effects of more width: feature formation (easier to instantiate additional task‑useful directions) versus packing efficiency (easier to spread existing directions). Early on, packing dominates; around $d\!\approx\!k$ capacity is spent on forming new directions (SI/WNO rise); once formation saturates, packing dominates again and SI/WNO drop. Centroids are tied directly to the BCE objective, so their feature formation saturates earlier, explaining the shorter first phase.
4.3 Topology‑induced superposition
On CONJUNCTION we train 3‑layer GCN, GIN, and GATv2 models (16 channels per layer), apply a global $\mathsf{mean}$ pool followed by a ReLU, and fit a two‑logit readout $(y_{A},y_{B})$. For each architecture we repeat training over 250 seeds. We use probes to study two feature families: (i) node‑level $\mathsf{Inside}_{\ell}(i)$ on the last pre‑pooling layer, which activates if node $i$ is part of an $\ell$-cycle, and (ii) graph‑level $\mathsf{Has}_{\ell}(G)$ on the post‑pooling representation, which activates if the graph $G$ contains any $\ell$-cycle.
@@ 4.3 Topology‑induced superposition
Despite $d\!\gg\!k$, superposition remains large. With k=4k{=}4 motifs and width d=16d{=}16, independent axes would yield $\text{SI}\!\approx\!1$ and $\text{WNO}\!\approx\!0$. Instead, Fig. 2 shows clusters at $\text{SI}\!>\!1$ and $\text{WNO}\!\gg\!0$, i.e., above random overlap. In most runs, at least one motif pair has $|\cos|>0.9$ (Table 1), most often $(C_{3},C_{4})$ or $(C_{5},C_{6})$; in rare cases ($\sim$1% in GIN) all four are nearly collinear, consistent with an emergent lever of ’cycleness’. High accuracy can still be achieved, suggesting that the pooling / reading takes advantage of residual multicoordinate structure or magnitude even when node‑level directions coalesce.
Cosine geometry reveals a cyclic structure. Figure 3 reports mean cosine similarity matrices. At the node level, the lengths of the nearby cycles align (for example, $C_{3}$ with $C_{4}$, $C_{5}$ with $C_{6}$) while distant pairs are weakly aligned or slightly antialigned. From the standard spectral view, a depth-$L$ message-passing GNN implements a low-degree polynomial filter in the normalized adjacency matrix $\mathbf{\tilde{A}}$ (Defferrard et al., 2016; Kipf and Welling, 2017): ignoring nonlinearities one can write
$\mathbf{H}^{(L)}\approx\sum_{k=0}^{L}\mathbf{\tilde{A}}^{k}\mathbf{X}\Theta_{k},$
where $\mathbf{X}\in\mathbb{R}^{n\times d_{\text{in}}}$ are the input node features, $\mathbf{H}^{(L)}\in\mathbb{R}^{n\times d_{\text{out}}}$ are the layer-$L$ node embeddings, and each $\Theta_{k}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ is a learnable coefficient matrix specifying how information arriving via $k$-step walks is mixed across feature channels. The power $\mathbf{\tilde{A}}^{k}$ encodes $k$-step walks on the graph, and its diagonal entries $(\mathbf{\tilde{A}}^{k})_{ii}$ correspond to closed walks of length $k$ at node $i$. Detectors for $C_{3},\dots,C_{6}$ therefore necessarily reuse many of the same short closed-walk monomials, which naturally makes nearby cycle lengths more aligned in representation space. After pooling, geometry re-mixes toward the task: the post-pool ReLU and linear readout learn graph-level axes that are linear combinations of node-level detectors (e.g., in GIN with mean-pool, $C_{3}$ aligns with $C_{6}$ while $C_{4}$ opposes $C_{5}$, which helps reject single-motif false positives).
Figure 3: Mean cosine similarities for the node features (top) and graph features (bottom) on Conjunction.
@@
Max pooling reduces lever sharing. Since a linear probe’s sign can flip with an equivalent boundary, we track $|\cos|$ to detect shared levers. Switching from mean to max pooling reduces node‑level alignment across all architectures: both the fraction of runs with any $|\cos|>0.9$ pair and with all pairs >0.9>0.9 drop (Table 1). Intuitively, a per-channel max cannot encode a conjunction on a single channel – one large activation would mask the other—so the model is pressured to place motifs on distinct channels, lowering $|\cos|$. The ordering in Table 1 (GCN >> GAT >> GIN in lever sharing) aligns with each layer’s mixing bias. GCN’s degree normalized sum acts as a low‑pass smoother, concentrating signals into a few graph‑harmonic axes and driving different motifs to share directions. GAT’s learned attention reduces indiscriminate mixing and yields moderate de‑alignment, but a single head offers limited selectivity. GIN’s unnormalized sum plus post aggregation MLP provides the most expressive channel rotation and gating, producing more specialized directions and the lowest node and graph level alignment.
Discussion. Even with generous width, semantically different topology features occupy similar node‑space directions. Max pooling alleviates but does not eliminate this, probably reflecting message-passing inductive biases. Note that many high‑accuracy models sit at $\text{SI}>1$ and $\text{WNO}\gg 0$ (Fig. 2), indicating that accurate solutions often rely on node‑level lever sharing with disentanglement deferred to the graph‑level readout – superposition here is not merely a symptom of under‑capacity.